OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half.

OpenStudy (anonymous):

A. y - 6 = one half(x + 3) B. y = one halfx - 6 C. y + 3 = one half(x - 6) D. x - 2y = 12

OpenStudy (anonymous):

2y = x -12

OpenStudy (anonymous):

That's not one of the answers, can you explain?

OpenStudy (anonymous):

y = x/2 - 6

OpenStudy (anonymous):

Again you aren't helping... could you explain the steps? I have no idea how to do this.

OpenStudy (texaschic101):

point slope form is y - y1 = m(x - x1) using slope(m) = 1/2 (6,-3) x1 = 6 and y1 = -3 now we just sub y - (-3) = 1/2(x - 6) = y + 3 = 1/2(x - 6)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Okay, thank you! I understand how you substituted the data, thank you! :)